Find the reference angle for -1060 degrees.
Found 2 solutions by math_tutor2020, greenestamps:
Answer by math_tutor2020(3835)


Answer: 20 degrees

Explanation

The angle -1060 is not between 0 and 360.

We need to add 360 to it repeatedly until we get inside this interval.

-1060+360 = -700
-700+360 = -340
-340+360 = 20

Or you can do this slight shortcut
-1060+3*360 = 20

Therefore, angles -1060 degrees and 20 degrees are coterminal.
They point in the same direction.

Because we ended up with a result in the interval 0 < x < 90, we don't need to do any more adjustments to get the reference angle.
This reference angle is in the first quadrant.
If we ended up in another quadrant, then we'd have to do a slight adjustment to it.

20 degrees is the final answer.

Answer by greenestamps(13334)



First, we need to add multiples of 360 degrees until the angle measure is positive. We might need to do more work after that.

Find the first multiple of 360 degrees that is greater than 1060:
360*1 = 360
360*2 = 720
360*3 = 1080

(-1060)+1080 = 20

The reference angle for any angle is the positive angle between the given angle and the x-axis. Since 20 degrees is in quadrant 1, the reference angle is 20 degrees; we don't need to do any more work.

ANSWER: 20 degrees